OpenStudy (anonymous):

For the circle x^2 -6x +y^2+8y +5=0 find the center, radius, and the equation of the tangent line at the point (1,-8)

OpenStudy (cwrw238):

x^2 -6x +y^2+8y +5=0 complete the square: ( x - 3)^2 - 9 + (y + 4)^2 - 16 + 5 = 0 ( x - 3)^2 + (y + 4)^2 = 20 compare this with the general form ( x - a)^2 + (y - b)^2 = r^2 where center is (a,b) and radius = r so from this u can see what center and radius are.

OpenStudy (cwrw238):

get it?

OpenStudy (anonymous):

So the center is 3,-4)?

OpenStudy (cwrw238):

thats right

OpenStudy (anonymous):

And is radical 20 the radius or

OpenStudy (cwrw238):

radical 20 is right

OpenStudy (anonymous):

Should radical 20 stay like that or would 2 radical 5 be a better answer

OpenStudy (cwrw238):

2 radical 5 is simplest form - yes

OpenStudy (cwrw238):

and how would you find equation of the tangent?

OpenStudy (anonymous):

No idea, I am more confused about that than I am about the center and radius.

OpenStudy (cwrw238):



OpenStudy (anonymous):

So you take your center point and the given point and find the slope and do the inverse of the slope the find slope of tangent? But how do you find the "b" variable of the tangent line

OpenStudy (cwrw238):

we know coordinates of center C and T is (1, -8) and the slope of the tangent at T has a slope -1/ slope of CT

OpenStudy (cwrw238):

use the general form y-y1 = m(x - x1) m = slope of tangent and (x1,y1) = (1,-8)

OpenStudy (anonymous):

Thanks you so much, I believe I understand how to do it now!

OpenStudy (cwrw238):

great - yw